### v1

@@ 4 Proposed method
DVFL splits the training into two phases coordinated by the owner. The first is an unsupervised and decoupled greedy guest-host representation learning phase. In this phase, the guest and host models learn feature extractors. The second is a supervised owner transfer-learning phase where the label owner learns to aggregate encodings from the host space and produce label predictions.
Within the guest-host phase, participants train their models asynchronously on their own obejctives. This allows some flexibility regarding when participants communicate and when they train. Since guests do not rely on feedback from hosts to update their models, host or communication faults are irrelevant to the guest models. Hosts store activations from guests as they arrive and use the latest message for their forward pass, so a missing input on the forward pass is no longer critical.
After hosts and guests complete training, the label owner trains its model. For each entity in $\mathcal{S}_{\mathrm{labels}}$, the owner requests guests and hosts to send its encoded and aggregated features. The owner uses this as an input for its task model, e.g. a classifier.
We present a DVFL approach to training split neural networks in the follwing subsections. The resultant algorithm is formalized in Algorithm 1. We also show how DVFL guests might exploit samples outside the joint sample space $\mathcal{S}_{\mathrm{guests}}$ to train better guest and host models.
@@ 4.1 The DVFL Hierarchy
Guests: The role of a guest in DVFL is identical to a VFL guest’s (see Section 2.1), i.e. guests $g_{i}\in\mathcal{G}$ are the agents $g_{i}$ who own datasets $\mathcal{D}_{i}$ with records of features $\mathcal{F}_{i}$ for the set of entities whose indices are present in $\mathcal{S}_{i}$. Each guest has a guest model and an unuspervised objective.
@@
    Input replay buffer $\mathbf{A}_{i}$. The host $h_{i}$ uses this to store a history of model inputs, if it wishes to reuse them.
Each host model has the same architecture but its own independent parameters $\theta_{h,i}$. It is possible that $\mathcal{H}\cap\mathcal{G}\neq\phi$.
Owner: The owner agent $o$ executes each training and communication round. It has access to ground truth labels, and is responsible for training a “final” or “owner” model on concatenated representations from members of $\mathcal{H}$ (i.e. it is a transfer learning model). The owner has the following:
  1. 1.
    Labels $L=\left\{\mathbf{y}_{j}\in\mathbb{R}^{\text{out}}:\text{label }\mathbf{y}_{j}\text{ of }\mathbf{x}_{j},j\in\mathcal{S}_{\mathrm{labels}}\right\}$:
    • $\mathcal{S}_{\mathrm{labels}}$ is the set of sample indices for which the host has the corresponding labels
    • We assume that the intersection of all the participants’ indices is not null to allow for training, i.e., $\left|\mathcal{S}_{\mathrm{guests}}\cap\mathcal{S}_{\mathrm{labels}}\right|>0$.
  2. 2.
    Final model $m_{o}(\text{ }\cdot\text{ };\theta_{o}):\mathbb{R}^{k_{H}\cdot{|\mathcal{H}|}}\to\mathbb{R}^{\text{out}}$ with parameters $\theta_{o}$.
The label owner may be a member of $H\cup G$. In any case, we refer to it shorthand as “the owner”. Figure 2 illiustrates a DVFL system with three guests and two hosts.
Figure 2: A DVFL system with three guests and two hosts. Guests train their local models on unsupervised objectives and hosts also train their aggregating models on unsupervised objectives. After that, the label owner trains a transfer learning model (such as a linear classifier head) on the encodings from the hosts.
@@ 4.2 Asynchronicity in DVFL
To avoid a single point of failure, host models should continue training in the absence of inputs from some guests. We adapt the use of activation replay buffers of Belilovsky et al. (2020) to facilitate training under guest and communication faults. We use single-reader single-writer “communication” registers so that hosts can reuse old activations from failed guests. A host reads the latest messages in its registers and uses them to fulfill its model inputs.
If hosts wish to train for more iterations than they receive activations for, they may store their model inputs in replay buffers. They may then reuse inputs as needed. We discuss these mechanisms in further detail in Appendix D.
@@ 4.4 DVFL for Split Training of Neural Networks
Given the DVFL hierarchy, ANNs are natural choices for participants’ models. Unsupervised neural networks are feature extractors Intrator (1992); Becker & Plumbley (1996), which is precisely the role of guest and host models. Similarly, the strong task performance of ANNs have been the primary driver of deep learning over the past few years, so they are suitable for owners, too. Indeed, just as SplitNN is a popular implementation of VFL, we expect ANNs to be common models for DVFL participants.
@@
Figure 1 illustrates the relationship between DVFL and VFL in the context of a global ANN model. We may say $m_{i}$s are all ANNs with parameters $\theta_{i}$, $m_{h,j}$ are also ANNs whose input is the concatenation of the outputs of all $m_{i}$s, and $m_{o}$ is an ANN whose input is the concatenation of the outputs of all $m_{h,j}$s.
Algorithm 1 DVFL training for split neural networks

### v2

@@ 5 Proposed Method
DVFL splits the training into two phases coordinated by the owner. The first is an unsupervised and decoupled greedy guest-host representation learning phase. In this phase, the guest and host models learn feature extractors on their own unsupervised objectives. The second is a supervised owner transfer-learning phase where the label owner learns to aggregate encodings from the host space and produce label predictions. We do not consider faults in owner training. They do not affect the guests and hosts which train the bulk of the network. Even if the owner were unstable, it only needs to collect hosts’ final encodings for each input once– then, it may train its model offline whenever it is up. [^2] This raises the question– why not do this between the guest and the host as it is in FedOnce (Wu et al., 2022)? Firstly, the owner’s job is only to predict, so a deep model is not required. As such, the dimensions of the embeddings and the computational cost of training its model will typically be small enough to neglect the cost of retrial, unlike a FedOnce/VFL host. Moreover, if the aggregator only gets one round of communication as in FedOnce, model performance is hindered. We explore this in our experiments and the Appendix.
Since guests do not rely on feedback from hosts to update their models, host or connection faults during the backward pass are irrelevant to the guest models. Hosts store activations from guests as they arrive and use the latest message for their forward pass, so a missing input on the forward pass is no longer critical.
The lack of end-to-end feedback combined with the three-tier hierarchy leads to significant robustness. The task of a host’s model is only to learn an encoding of features in the joint guest domain, which is not an instance-specific task. Therefore, a host may still update its model meaningfully with some out-of-date activations. Similarly, guests may use their entire datasets as opposed to only members of the intersection. Besides leading to better models, using more data means a single fault does not affect the overall uptime as much. On the backward pass, faults only affect the crashed participant.
@@ 5.1 The DVFL Hierarchy
Guests: The role of a guest in DVFL is identical to a VFL guest’s (see §2.2), i.e. guests $g_{i}\in\mathcal{G}$ are the agents $g_{i}$ who own datasets $\mathcal{D}_{i}$ with records of features $\mathcal{F}_{i}$ for the set of entities whose indices are present in $\mathcal{S}_{i}$. The guest model’s role is still to encode its dataset’s private features. Unlike VFL, they learn to do so on local unsupervised objectives.
@@
    Input replay buffer $\mathbf{A}_{i}$. The host $h_{i}$ uses it to store a history of model inputs if it wishes to reuse them.
Each host model has the same architecture but its own independent parameters $\theta_{h,i}$. Unlike standard VFL, the hosts do not make predictions. Their models are only responsible for aggregating guest outputs and learning to effectively map them to a lower dimensional space. Note that when there is no such compression, the host can achieve this without any learnable parameters– only a single aggregating layer e.g. concatenation is necessary.
Owner: The owner agent $o$ executes each training and communication round. It has access to ground truth labels, and is responsible for training a “final” or “owner” model $m_{o}(\text{ }\cdot\text{ };\theta_{o}):\mathbb{R}^{k_{h,1}}\times\dots R^{k_{h,|\mathcal{H}|}}\to\mathbb{R}^{\text{out}}$ on representations from members of $\mathcal{H}$ (i.e. it is a transfer learning model). The owner has access to the ground truth targets $L=\left\{\mathbf{y}_{j}\in\mathbb{R}^{\text{out}}:\text{label }\mathbf{y}_{j}\text{ of }\mathbf{x}_{j},j\in\mathcal{S}_{\mathrm{labels}}\right\}$.
Figure 2: A DVFL system with three guests (a) and two hosts (b). Guests train their local models on unsupervised objectives and hosts also train their aggregating models on unsupervised objectives. After that, the label owner (c) trains a transfer learning model (such as a linear classifier head) on the encodings from the hosts.
@@ 5.2 Asynchronicity in DVFL
To avoid a single point of failure, host models should continue training without inputs from some guests. Single-reader single-writer “communication” registers allow hosts to reuse old activations under guest and communication faults. Hosts can store model input history in replay buffers for reuse if they wish to train for more iterations than they receive activations for. We discuss these mechanisms in further detail in Appendix F.
@@ 5.4 DVFL for Split Training of Neural Networks
NNs are natural choices for participants’ models. Unsupervised neural networks are feature extractors (Intrator, 1992; Becker & Plumbley, 1996), which is precisely the role of guest and host models. Similarly, the strong task performance of NNs have been the primary driver of deep learning over the past few years, so they are suitable for owners, too. Indeed, just as SplitNN is a popular implementation of VFL, we expect NNs to be common models for DVFL participants.
@@
Figure 1 illustrates the relationship between DVFL and VFL in the context of a global NN model. We may say $m_{i}$s are all NNs with parameters $\theta_{i}$, $m_{h,j}$ are also NNs whose input is the concatenation of the outputs of all $m_{i}$s, and $m_{o}$ is an NN whose input is the concatenation of the outputs of all $m_{h,j}$s. The pseudocode is presented in the Appendix A (Algorithm 1).
6 Experiments
Datasets
As a toy example, we consider a system of $|\mathcal{G}|$ guests which can each see $\frac{1}{|\mathcal{G}|}^{th}$ of an handwritten digit from the MNIST dataset (Deng, 2012). For our experiments, we set $|\mathcal{G}|=4$ i.e. $g_{1}$’s dataset contains the top $28\times 7$ pixel patch of an MNIST image, $g_{2}$’s dataset contains the second , $g_{3}$’s contains the third, and $g_{4}$’s contains the bottom $28\times 7$ patch (see Fig. 3(a)). We also experiment on some benchmark VFL datasets in the Appendix.
New Baselines and Loss Functions
DVFL is an alternate strategy to VFL. Naturally we compare DVFL for NNs to SplitNN, a direct application of VFL to NNs. Naive VFL implicitly employs the “wait” protocol discussed in §2.4.1.
We also try the modified timeout strategy from §4 by imputing missing inputs with zeros (zeros), stale activations (buffer), and skipping that round altogether (skip). For SplitNN hosts and DVFL owners, we use cross-entropy loss. For DVFL hosts and guests, We use a learnable decoder head and MSE reconstruction as an unsupervised objective.
FedOnce Baseline
We also provide a baseline with FedOnce (Wu et al., 2022). FedOnce is similar to DVFL in the sense that guests train their own models on unsupervised objectives, and similar to VFL in the sense that there is a centralized aggregating host that owns labels. FedOnce only requires one round of communication (the guest-encoded dataset after all guests have fully trained). It can be argued that each guest can retry until the host has heard from all of them (since it would not be holding up training of any future iterations). If that is implemented, faults are not a worry. Regardless, we also try an alternate fault-handling strategy inspired by VFL-zeros for compatibility with deadline-based scheduling.
Hyperparameters, model details and ablation study are presented in the Appendix.
Asynchronous Baselines
We also compare DVFL to the two discussed viable asynchronous VFL algorithms: VAFL (Chen et al., 2020a) and VFB2 (Zhang et al., 2021). For VAFL, we eschew random batch sampling for iteration (in the same order by all guests). This locality gives us a skyline of VAFL’s performance– non-active guests’ “stale” activations will be recent. For VFB2, we simulate without the two distinct communication graphs to minimize the scope for faults. Further, since we experiment with only one label-owning party, the “active parties” class only contains a single host ($m=k=1$ by the convention of the original paper). Also, since the aggregation strategy is summation, we modify the first layer of the host model’s input dimension to equal a single host’s output dimension.
FedVS Baseline and More Datasets
We also compare the model performance of SplitNN and DVFL to results cited in the work FedVS (Li et al., 2023). Since FedVS is incompatible with our fault model, we only compare performance under perfect conditions. We use the same datasets (Yeh & hui Lien (2009); Markelle Kelly ; Li et al. (2022); Sakar et al. (2019)) as used in the experiments for FedVS. FedVS restricts its models to use polynomial networks (Livni et al., 2014a) i.e. equally-sized linear layers with polynomial activation functions. The models we use for SplitNN and DVFL are smaller– each layer is half as wide as the next, and we use Leaky ReLUs.
6.1 Fault Simulation
Six hyperparameters control the fault model described in §2.4.1. $R^{(d)}_{i}$ is the fault rate for a live $i\in\{\mathrm{connection},\mathrm{guest},\mathrm{host}\}$. $R^{(u)}_{i}$ is the chance a dead $i$ rejoins the network before the next request. We discuss how exactly the faults are simulated in the Appendix. We assume no faults during label supervision or inference.
6.2 Experimental Setup
Table 1: We compare DVFL to standard VFL. After filtering outliers by IQR, we report the mean and 2x the standard deviation of MNIST test performance when trained with various entities $i\in\{\mathrm{connection,guest,host}\}$ susceptible to fail at a rate of $R^{(d)}_{i}=0.3$. A failed resource has $R^{(u)}_{i}=\mathbb{P}_{\mathrm{rejoin}}\in\{1,0.5,0.1\}$ to be available next time it is requested. SplitNN, the baseline VFL algorithm, cannot train at all when there are any faults. On the other hand, our DVFL implementation degrades gracefully with faults. Using an explicit fault handling strategy within VFL does allow it to train. But the degradation with faults is sharper than DVFL and the variance in model performance is higher. Under fault-free circumstances, SplitNN slightly outperforms our DVFL implementation. With minor adjustments, SplitNN can tolerate some faults, but our approach is more resilient and stable.
Strategy Test Accuracy (%) with 0.3 Loss Rate
No Faults Connection Loss Guest Loss Host Loss
DVFL-NN 97.80$\pm$0.12 97.78$\pm$0.11 97.78$\pm$0.14 97.72$\pm$0.17 97.77$\pm$0.11 97.74$\pm$0.16 97.58$\pm$0.20 97.81$\pm$0.12 97.79$\pm$0.14 97.60$\pm$0.33
VFL (SplitNN) 97.85$\pm$0.17 N/A N/A N/A N/A N/A N/A N/A N/A N/A
VFL (SplitNN-skip) 97.86$\pm$0.16 97.50$\pm$0.54 96.60$\pm$0.51 77.21$\pm$5.34 97.44$\pm$0.57 96.65$\pm$0.45 80.22$\pm$4.31 97.65$\pm$0.41 97.59$\pm$0.50 97.29$\pm$0.89
VFL (SplitNN-zeros) 97.84$\pm$0.15 97.64$\pm$0.38 97.53$\pm$0.40 97.44$\pm$0.36 97.68$\pm$0.40 97.59$\pm$0.40 96.95$\pm$0.50 97.68$\pm$0.30 97.65$\pm$0.37 97.31$\pm$0.93
VFL (SplitNN-buffer) 97.85$\pm$0.17 97.66$\pm$0.38 96.85$\pm$0.74 93.45±$\pm 2.10$ 97.59$\pm$0.58 96.76$\pm$0.70 91.78$\pm$1.70 97.81$\pm$0.32 97.71$\pm$0.40 97.69$\pm$0.51
VFL (FedOnce-zeros) 97.76$\pm$0.20 96.89$\pm$0.29 96.41$\pm$0.30 92.80$\pm$0.75 97.47$\pm$0.37 97.45$\pm$0.31 97.50$\pm$0.31 97.28$\pm$0.35 97.34$\pm$0.37 96.93$\pm$0.51
VAFL (skyline) 97.57$\pm$0.30 97.17$\pm$0.46 96.88$\pm$0.36 96.46$\pm$0.44 97.18$\pm$0.42 96.99$\pm$0.44 96.45$\pm$0.51 97.17$\pm$0.48 96.89$\pm$0.36 96.42$\pm$0.52
VFB2 (skyline) 97.72$\pm$0.16 97.43$\pm$0.38 97.33$\pm$0.41 96.09$\pm$0.47 97.36$\pm$0.45 97.24$\pm$0.41 96.29$\pm$0.42 97.43$\pm$0.49 97.20$\pm$0.61 96.79$\pm$0.90
Performance under Faults
In order to show that it is possible to use DVFL even with crash faults, we train DVFL with $|\mathcal{H}|=4$ hosts with a fault rate of 0.3 and a rejoin rate of 0.1. For a given experiment, the failure mode (connection, guest, or host) is fixed.
Effect of Redundancy
In order to show that redundancy of hosts improves model performance and fault tolerance, we simulate DVFL training with $|\mathcal{H}|$ = 1, 2, 3, and 4 hosts. We measure test accuracy given connection fault rate $\in\{0,0.3,0.6\}$ during training.
Learning from a Limited Intersection
We first fix a small labelled and aligned intersection size. The remaining sample space is shuffled and distributed evenly across all four guests. For example, if we wish to use $1024$ labeled samples, we shuffle samples indexed $1024-59999$ and split the $58976$ samples evenly, such that each guest gets $14744$ unique MNIST patches + $1024$ intersecting MNIST patches = $15768$ total samples. We try this experiment with $128$, $256$, $512$ and $1024$ labeled samples. We compare with SplitNN trained on only the labeled samples.
7 Results and Discussion
(a) For our experiments, the input received by each guest for a given sample, as well as the reconstruction its autoencoder model produces.
[…]
(b) Model performance improves with redundancy, especially when there is a nonzero crash rate.
7.1 Graceful Degradation with Faults
(Table 8) shows that model performance degrades with faults, but not catastrophically (unlike VFL with SplitNN). In ideal conditions, DVFL performance is slightly poorer than SplitN’s. This is not unexpected as we will discuss in §9 Observe that the degradation with faults in DVFL is much more graceful and stable than the SplitNN variants.
In the case of SplitNN-based algorithms, a connection and guest fault are functionally identical. In the case of DVFL, there are some performance savings under connection faults– the corresponding guest and host can still meaningfully learn even in the absence of connection. Further, redundant hosts may continue training as normal.
Tolerating some faults pays off
The artifical input induces a Dropout effect (Srivastava et al., 2014). This reduces bias but competes with information loss in the model. DVFL models, without cross-device feedback, see smaller information loss compared to SplitNN-based methods (which do not respond to faults as well, but still exhibit somewhat graceful test accuracy degradation). The simplest SplitNN-based method, Skip, leads to the worst performance since information from non-faulty guests is also lost. Zeros works as if failed neurons do not exist at all in simple linear and convolutional layers. This would fail with more complex architectures and algorithms e.g. skip connections and constrained solvers. Buffer empirically cannot exploit locality in activations and performs poorly. Reconstructing missing inputs based on statistics, as Fu et al. (2022b) do for gradients, may be the best solution but would suffer from bias and computational infeasibility.
Comparison to asynchronous VFL
Asynchronous algorithms suffer from higher variance, perhaps since the information loss from a single failure do not have the safety net of information from successful interactions. Further, these skylines perform significantly worse than VFL and DVFL even when there are no faults. In fact, we argue that VFL-zeros provides a skyline for performance from any asynchronous algorithm (at least when the model has no skip connections or optimization constraints). As discussed in the previous paragraph, VFL-zeros acts as if failed neurons did not exist at all. This is as if an oracle told an asynchronous host which guests/connections will be live, and the host calculated losses based on multiple guests’ embeddings at once hence leading to lower-variance loss estimates. VFB2 and other algorithms using summation-based aggregation also suffer from constraints on the model architecture. The output of each guest must have the same dimension. Further, the host input dimension is smaller than what it would be if it were to aggregate by concatenation, so there are fewer degrees of freedom for the model.
Comparison to FedVS and results on other datasets
FedVS (Li et al., 2023) provides both gradient privacy (via cryptographic multi-party communication) and straggler resistance. The information loss thanks to quantization and random masking in FedVS outweighs that from localized training in DVFL. As a result, performance lags even in perfect conditions (Table 2). We also see that DVFL performance tracks SplitNN’s closely on all datasets, including tabular ones which are typically “hard” for neural networks.
Table 2: On a variety of typical vertically partitioned datasets, we see that DVFL performs comparably to SplitNN even under perfect conditions. Even on tabular datasets harder (especially for NNs) than MNIST such as Credit Card and Parkinson’s, DVFL model performance is comparable to SplitNN. Despite using a smaller model, both outperform the results cited in the FedVS paper (Li et al., 2023).
Strategy Test Accuracy (%) Desiredata
Dataset Privacy Fault Tolerance
Credit Card Parkinson’s CalTech-7 Handwritten
DVFL 82.66 85.76 99.00 98.25 No Gradients Crash
VFL (SplitNN) 84.11 86.64 99.66 98.25 None None
FedVS (Li et al., 2023) [^3] Results extracted from graphs. Raw figures were unavailable and the authors did not respond to our inquiries. 81.99 81.83 95.98 98.25 Quantization and Noise Straggler
005555101010101515151520202020252525259797979797.297.297.297.297.497.497.497.497.697.697.697.697.897.897.897.8Guest training epochs per communication epochMNIST Test Accuracy (%)
Comparison to FedOnce
FedOnce (Wu et al., 2022) aims to reduce the volume of communication at the cost of model performance. Equivalently, Belilovsky et al. (2020); Siddiqui et al. (2023) demonstrate that updating all layers on local objectives on each forward pass performs better than fully training and freezing a layer before moving on to the next. Even if we consider this completely fault tolerant, the overall performance loss from reduced communication is enough that DVFL performs better or comparably even in our experiments with the high fault rates. If communication size is a concern, DVFL provides fine-grained control over that. This is discussed further in §8.1.
7.2 Redundancy Improves Performance, Asynchronicity Affords Efficiency
We observe (Figure 3(b)) that test accuracy improves with the number of hosts with all other system parameters fixed. This is the case even when there are no faults. This is not surprising, as the owner’s aggregation of host results can be seen as a form of bagging (Breiman, 1996). As a corollary, the degradation of performance with connection faults is more graceful with increasing redundancy.
7.3 Out-of-Intersection Training and Privacy
Although VFL with SplitNN marginally outperforms our DVFL NN in the experiments with perfect conditions, DVFL significantly outperforms VFL on a limited labeled intersection. DVFL’s advantage is greater with fewer labels (Table 3). DVFL is performant while also completely eliminating the possibility of gradient-based inference attacks. As reviewed in 3, complex mechanisms can be developed on top of SplitNN to provide privacy guarantees within some bounds. DVFL achieves complete privacy against gradient-based attacks without any extra computation and minimal performance loss.
Strategy Test Accuracy (%)
# of Labeled Samples
128 256 512 1024
DVFL 76.45 82.51 85.18 88.90
VFL (SplitNN) 65.07 77.51 83.09 87.26
8 Engineering Flexibility with DVFL
8.1 Control over Communication Volume
Since guests and hosts both train on their own objectives, it is not necessary that they synchronize on every iteration. This means communication cost can be controlled by only transferring some embeddings, at the cost of performance. We experiment with various communication frequencies and present the results in Table 3. These experiments are discussed in further detail in the Appendix F.1.
8.2 Control over Computational Cost
Ostensibly, DVFL is slightly more computationally expensive than VFL since both guests and hosts must generate their own update signals in the absence of supervisory labels. In our experiments with autoencoders, this manifests as training the decoder head. However, DVFL allows the flexibility of allowing each participant to train for a different number of iterations. This is unlike VFL where the guest must compute a forward pass and then update every time the host makes a request. As such, the number of iterations for each participant can be tuned to balance model performance and computational cost.
Our main experiments with DVFL only use 20 epochs of guest training while VFL hosts, DVFL hosts and DVFL owners use 60 epochs. We measure the wall-clock time of each process in the Appendix F.3 and see that the guest processes are actually alive for a shorter duration in DVFL in comparison to VFL ($\approx 55.88$s vs $\approx 80.22$s under the same operating conditions). Indeed, the same way that the meaningful learning is possible even if the guest sends a limited number of activations, the host may make do with storing a limited number of activations. In the Appendix F.2 we discuss how a memory constrained host may still meaningfully learn while only storing the last few activations rather than the entire history.
9 Limitations
Our models are shallow. Wang et al. (2021) show that the performance improves with the length of backpropagation in blockwise-trained networks, and suggest that it is due to increasing mutual information between the shallower and deeper layers. Further, given a model, the loss does not guide the model to extract strictly task-useful features. These issues typically also lead to scalability issues to harder datasets, although we observe in our experiments on tabular datasets that the difference is not that large. However, we expect that the information loss from the absence of direct connections between guests’ neurons is greater than that due to the loss of task-relevant information.
10 Broader Impact
We believe DVFL will enable large scale and democratic participation in joint learning tasks where VFL has been unable to provide solutions. e.g. street & private residential cameras predicting traffic incidents at a busy intersection. Participants might not all be able to see each training incident i.e. small $\mathcal{S}_{\text{guest}}\cap S_{\text{labels}}$, or might be concerned about privacy, or suffer from intermittent connection. DVFL is also more practical for existing cross-silo applications. Consider the typical example of various specialized healthcare providers predicting, say, cancer. Not only can the hospitals train asynchronously and not worry about faults, they need not worry about compromised hospitals launching gradient inference attacks. Further, they may now admit more hospitals to their system even if they have few patients in common. We see DVFL as a step towards usable and ubiquitous machine learning.
11 Conclusion and Future Work
We present Decoupled Vertical Federated Learning, a strategy for joint learning on vertically partitioned data. Instead of adding complexity to cope with the data leakage and synchronization associated with BP, we eschew feedback altogether. Combined with the ability to reuse feedforward signals, this allows for fault tolerance in training.
All-in-all, we address three impracticalities with VFL training: single points of failure, training data being limited to the intersection, and gradient-based privacy attacks: Training may continue even when there are faults, or there are a limited number of samples common to all guests. Malicious or compromised guests can no longer use gradient feedback to figure out possibly sensitive information. Existing techniques to enforce privacy guarantees either hamper model convergence, are expensive to implement or both.
In turn, DVFL introduces new flexibility and practicality to systems intended for learning from vertically partitioned data– the volume of communication, memory usage and overall complexity can be controlled by a designer.
Our results suggest DVFL’s performance to be comparable to SplitNN on various ML tasks. Unlike SplitNN, DVFL degrades gracefully with manykinds of loss faults. Even with explicit fault handling strategies, SplitNN’s performance drops off more sharply than DVFL’s. In general, DVFL’s asynchronicity provides a great deal of engineering flexibility. Some of it is discussed in the Appendix.
The next step for DVFL is to generalize it for more tasks and general architectures. VFL is readily compatible with a variety of machine learning models (including ensemble models). With a judicious choice of training objective for guest models, DVFL may also be extended past NNs.
Appendix A DVFL for Split Neural Network Training
Algorithm 1 DVFL training for split neural networks (DVFL-NN)